OpenStudy (anonymous):

Factor completely: -5k2 + 20k - 30

OpenStudy (anonymous):

-5k^2 + 20k - 30=(20-5 k) k-30=-5 ((k-4) k+6)=-5 (k-2)^2-10

OpenStudy (anonymous):

\[-5 (-i k+\sqrt(2)+2 i) (i k+\sqrt(2)-2 i)\]